The greater of two consecutive even integers is 6 less than three times the lesser. Find the integer

Answer by CubeyThePenguin(3113)

consecutive even integers: x, (x+2)

x + 2 = 3x - 6
8 = 2x
x = 4


greater integer = x+2 = 6